A company manufactures advance chips. For any random selected chip, the defective rate is 0.24. Find the probability that exactly 2 chips are defective in a batch of 12.
Answer #1

Given that ,

p = 0.24

1 - p = 1 - 0.24 = 0.76

n =12

Using binomial probability formula ,

P(X = x) = (n C x) * px * (1 - p)n - x

P(X = 2) = (12 C 2) * 0.242 * (0.76)n - x

probability =0.2444